OpenStudy (anonymous):

Find the surface area of the a cone which has dimensions of 7 and 3.

Directrix (directrix):

The surface area of a right circular cone is found by the formula: SA = pi*r^2 + pi*r*L where r is the base radius and L is the slant height. pi*r^2 is the area of the circular base.

Directrix (directrix):

SA = pi*r^2 + pi*r*L SA = pi* 4^2 + pi*4*7 SA - ?

Directrix (directrix):

@SadieK Note that the drawing shows the length of the base diameter, not its radius.

OpenStudy (anonymous):

is it 138?

Directrix (directrix):

Is that what you got?

OpenStudy (anonymous):

Yes.

Directrix (directrix):

I had 44 pi which approximates to 138.26.

OpenStudy (anonymous):

Thank you. I am going to close this question and ask my last one.

Directrix (directrix):

OK.